You throw a die repeatedly until you get a 6. What is the probability that you need
to throw more than 20 times to get 6?
1
Expert's answer
2020-04-28T18:35:26-0400

Throwing a die more than 20 times to get 6 means that 20 times you do not get 6. The probability of not getting 6 once is "\frac{5}{6}". The probability of not getting 6 during 20 times is "(\frac{5}{6})^{20} = \frac{95367431640625}{3656158440062976}" which is the answer.